### v1

@@ 4.3 Histograms and Log-normality
Similarly the expression for binormal distribution composed of two sub-populations having means $\mu_{1}$ and $\mu_{2}$ and standard deviation $\sigma_{1}$ and $\sigma_{2}$ is
@@
where $m_{1}$, $m_{2}$ and $s_{1}$, $s_{2}$ are location and scale parameters for the two superposed lognormal distributions in the bimodal lognormal distribution. Also, the fit parameters obtained from fitting the normal and lognormal distributions are shown in Table 3.
@@
We found that seven observations (with IDs 0080940101, 0124930501 0158961301, 0411780301 and 0411780301) out of 22 observations exhibit a bimodal distribution. In Flux Figure 4, we present the histogram for one of the observations which shows a clear bimodal distribution. Blazar flux distributions showing bimodal feature are also reported in H. E. S. S. Collaboration et al. (2010); Mohorian et al. (2022); Shah et al. (2018). The presence of the bimodal distribution suggests a clear distinction between low and high flux states. However, since the light curves often exhibit continuous flux variation, it is difficult to categorize specific flux points as belonging to either a quiescent state or a flaring state using a bimodal PDF.
@@ 4.4 Power spectral density analysis
Studying variability in AGN using methods based on both the time domain and frequency domain can help uncover its underlying physical processes. The Power Spectral Density (PSD) serves as an important analytical tool in the frequency domain, measuring the ’variability power’ at a given temporal frequency (or timescale). This enables us to understand the nature of variability, providing clues about its possible origin.The Discrete Fourier Transform (DFT) is widely used as a measure of PSD and is suitable for discrete set of observations.(see Vaughan et al., 2003; González-Martín & Vaughan, 2012; Mohorian et al., 2022) and at discrete times $t_{j}$ with $j=1,2,...,n$, which occurs at frequencies $\nu_{\rm min}=\frac{1}{T}$, $2\nu_{\rm min}$ , $\nu_{\rm max}=\frac{1}{2\Delta t}$, it can be expressed as,
@@
In this study, we performed the fitting of the PSDs using scipy curve-fit python package [^3] https://docs.scipy.org/doc/scipy/reference/generated/scipy.optimize.curve_fit.html. Our analysis of the XMM- Newton X-ray light curve of PKS 2155-304 reveals a predominance of red noise, which is well characterized by a PL model. The PSD slopes ($\beta_{\mathrm{P}}$), as illustrated in Figure 6 and quantified in column 8 of Table 2, exhibit significant variability across the 22 observations of PKS 2155-304. The mean $\beta_{\mathrm{P}}$ value is 2.49, with a range spanning from 0.922 (observation 0411780701) to 4.491 (observation 0411780201). These variable PSD slopes correspond to different stochastic processes: flicker noise ($\beta_{\mathrm{P}}\sim 1$), red noise ($\beta_{\mathrm{P}}\sim 2$), and other noise processes with steeper PL index, that is,   $\beta_{\mathrm{P}}>2$ (Press, 1978).
@@ 4.5.1 HARDNESS RATIO
It is interesting to note that although we observed pronounced flux variability throughout the observations, spectral variations measured by HR exhibit a complex pattern, with no significant variation in the majority of observations. Obs. IDs 0080940101, 0124930501, and 0411780501 reveal a trend (or a weak trend) of harder-when-brighter, whereas Obs. ID 0411780201 shows a trend of softer-when-brighter (e. g., see Bhatta et al., 2018). In many observations, the error in HR at lower count rates is also higher, thus not providing conclusive HR variations with count rate. Additionally, Figure 8 presents the distribution of HR across the entire set of observations as a function of count rate. Although the data reveal no simple correlation between flux and HR, the figure demonstrates the long-term spectral variability of the source across a wide range of flux states. Note that different observational epochs are plotted in distinct colors.
@@ 4.5.2 SPECTRAL FITTING
We also performed the spectral analysis of X-ray emission from blazar PKS 2155―304 by extracting the spectra of the observations within the energy range 0.3–10 keV and consequently, fitted them with PL models from XSPEC [^4] https://heasarc.gsfc.nasa.gov/xanadu/xspec/manual/XspecManual.html. To accurately characterize the underlying spectral shape of the emission, we employed the following three different PL models.
(i) Powerlaw (PL): This is a simple powerlaw model in which the photon count rate distribution over the energy range is expressed solely by a single power index. The model can be written as
@@
(ii) Broken powerlaw (BPL): In a BPL model, a spectral shape is characterized by two powerlaw indices meeting at a discontinuity at an energy point known as the break energy. The distribution can be expressed as:
$\dfrac{dN}{dE}=\cases{K}\cdot E^{-\Gamma_{1}}&\text{if $E\geq E_{b}$},\\ K\cdot E^{-\Gamma_{2}}\text{otherwise},{}$ (18)
Here $K$, $\Gamma_{1}$, $\Gamma_{2}$ and $E_{b}$ are normalization constant, low and high energy photon indexes and break energy, respectively.
(iii) Log Parabola (LP): Log parabola is one of the models widely used in the spectrum analysis of blazars (Massaro et al., 2004, 2006). The distribution function for LP is
@@
In this study, we fit the X-ray spectra obtained from XMM- Newton observations listed in Table 1 using three models available in the X-ray fitting package, xspec. To minimize instrumental artifacts, we restricted our analysis to the 0.3–10 keV energy range. We initially applied the simple PL model, which failed to yield a significant reduced chi-square ($\chi^{2}_{r}$). To improve the fit, we then applied more complex models: a BPL and LP. Both these models resulted in significantly improved $\chi^{2}_{r}$ values. Furthermore, we noted spectral curvature at lower energies. To accurately capture this feature, we utilized the LP model, a variant of the standard log-parabola model described in Equation 19. The results from our spectral fits are presented in Table 4, and a representative fitted spectrum from one observation is shown in Figure 9. The spectral fits for the remaining observations are presented in the Appendix. Our result indicate that BPL and LP model provides a good fit for all observations. A similar investigation of the spectral curvature of the PKS 2155-304 source was conducted by Gaur et al. (2017), utilizing XMM- Newton data from 2009 to 2014 and employing log-parabola model fitting techniques.
Table 4: Spectral properties of PKS 2155-304. Col. 2: observation ID; Col. 3: Spectral models, powerlaw (PL), broken powerlaw (BPL); Col 4: photon index (PL), high-energy photon index (BPL); Col. 5: curvature parameter (LP), low-energy photon index (BPL); Col. 6: break/peak energy in keV; Col. 7: $\chi^{2}$/degrees of freedom; .
@@ 5.1 Timing Analysis
Variability: The source displays substantial variability both on intra-day and long-timescale. As indicated in Table 2, on intra-day timescales, it exhibits an average fractional variability of $\sim 6\%$, with one of the observations showing variability as high as 20%. During the entire observational period, the fractional variability reaches as high as $\sim 47\%$. To put it in a perspective, the variability exhibits significant modulation with maximum flux levels approximately eight times higher than the minimum observed values.
The X-ray emissions of TeV blazars exhibit significant variability at intra-day timescales. This could possibly be attributed to the fact that the observed X-ray emission in HBLs is associated with synchrotron emission by a population of electrons located at the high-energy tail within the synchrotron component of their SEDs. The intra-day variability can be explained in the scenario of the turbulent jet of the blazar. In this framework, the observed variability can be interpreted as a superposition of fluctuations arising from numerous small-scale inhomogeneous components with varying Doppler boosting. Furthermore, as shocks propagate through the turbulent jet, energy dissipates from the largest to the smallest cells. Emission from some of these dominant smaller cells can contribute to short timescale variability. (Calafut & Wiita, 2015; Marscher, 2014; Bhatta et al., 2013). Using the variability timescale and the causality argument, the spatial scale of the inhomogeneity in the jet can be related to the size of the turbulent cells. An upper limit estimate for size of the compact spherical cells at redshift ($z$), corresponding to the minimum variability timescale ($\tau_{\rm var}$) and Doppler factor ($\delta$), can be computed using the following relation:
$R\leq\frac{\delta}{(1+z)}\cdot{c\tau_{\rm var}}.$ (20)
The lower limit of the Doppler factor associated with the source is of the order $\delta\sim$10 (Falomo et al., 2014; Kapanadze et al., 2020), which corresponds to a size of the emitting central region $R\sim 10^{14}$ cm using an average of minimum variability timescale over all the observations 2.0 ks. Similar results using Suzaku and NuSTAR observations are discussed in (Bhatta et al., 2018; Zhang et al., 2021). Furthermore, the variability timescales can be linked with the characteristics time scales such as cooling timescales of the source. In this particular case, the synchrotron cooling timescale in the observer’s frame is
@@
where $\delta$, $\gamma$, $B$ and $z$ represents bulk Doppler factor, electron Lorentz factor, magnetic field in Gauss and redshift parameter, respectively. An average cooling timescale of 2 ks, in combination with a magnetic field strength of B = 1 Gauss, typical for the jet’s magnetic field, results in an average electron energy of 4.32 × 104 Lorentz factors, enabling X-ray emission in the XMM- Newton band.
Seven out of 22 countrate distributions show bimodal nature while the remaining show a unimodal. During the study of flux distribution using long-term optical and gamma-ray observations, it was found that both the light curves being highly correlated both of their corresponding flux distribution was best represented by a log-normal probability density function (PDF) (Bhatta, 2021; Bhatta & Dhital, 2020). Such distributions are often interpreted within the framework of non-linear and multiplicative processes in the source. Interestingly, unlike optical and gamma-ray emission from the source, the overall X-ray flux histogram is very complex showing multiple peaks and therefore the distribution does not follow a similar PDF. It may be due to the fact that Optical and Fermi/LAT gamma-ray observations lie nearly at the synchrotron and high energy peak, respectively. However the X-ray emission lie near the tail of the synchrotron component and therefore explores that has highest energy electron population. A lack of any simple PDF indicate the multi-component X-ray emission of the sources.
@@
Our analysis reveals that the observed PSD slopes ($\beta_{P}$) for PKS 2155-304 predominantly fall within the red noise regime $(2\leq\beta_{P}\leq 4)$, with occasional instances approaching pink noise $(\beta_{P}\approx 1)$. These variable PSD slopes might suggest distinct local manifestations of stochastic processes occurring within the turbulent jet structure. However, it is also possible that the long-term light curve is actually characterized by a single flatter PL PSD. In this case, the intra-day observations represent only a small segment observed across different flux states, resulting in apparent variations in the local PSD slopes. However, we did not estimate the long-term PSD by combining all the observations, as such a PSD would be affected by the large gaps between observations spanning years. This would result in an inaccurate characterization of the long-term X-ray variability in the source
@@
The linear RMS-flux relation indicates that the variability properties of AGNs are closely linked across different flux states. This suggests that multiplicative non-linear stochastic process may be driving the observed variability, leading to a flux distribution skewed towards higher flux levels, such as a lognormal probability density function (PDF) (Uttley et al., 2005). Conventionally, the linear RMS-flux relation along with the lognormal flux distribution is linked to uncorrelated fluctuations in the $\alpha$-parameter in the accretion disk. These fluctuations are governed by viscosity changes occurring at various radii, which then propagate outward, gradually modulating the mass accretion rates on larger scales. (Lyubarskii, 1997). These disk modulations can propagate through the jet due to a disk-jet connection in radio-loud AGNs. In blazars, these modulations are further amplified by relativistic beaming effects, which may also be influenced by projection effects. More recently, (Kundu et al., 2022) presented an explanation for the RMS-flux based on the internal shocks in the jet scenario. In the simulation of the multi-frequency light curve from the jet, the relationship was observed when applying shocks with varying amplitudes, determined by the speed of the colliding blobs that produce in the shocks.
@@ 5.2 Spectral Analysis
In general, AGN spectra may be composed of various emission components, such as soft X-ray excess, neutral iron emission lines around 6.4 keV, and parts of either a synchrotron hump or a Compton hump. However, blazar spectral emission is primarily jet-dominated due to beamed electromagnetic radiation and the non-thermal emission in blazars usually follows a PL (Aharonian et al., 2005b). Studying the X-ray spectra of blazars helps us understand the particle acceleration mechanisms at play, the properties of the jet environment, and the overall emission processes that power these luminous objects. The excellent sensitivity of XMM- Newton provides detailed information about the emission mechanisms, which we obtain by fitting the X-ray spectra of the source.
@@
HR analysis reveals a moderate variation in fractional variability along with the variability amplitude in the entire observation period. Only two observations show a significant HR variation. In addition, we fitted the source spectra using three widely used XSPEC models: powerlaw, broken powerlaw, and log-parabola. These models are chosen to represent non-thermal emission from relativistic electrons in the magnetic field of the blazar jet. In the synchrotron emission scenario, a simple PL spectrum can be interpreted as coherent synchrotron emission produced by a simple PL distribution of particles. However, BPL and log-parabolic models suggest more complex scenarios. A BPL might indicate that the observed emission is contributed by multiple populations of particles with different distributions. On the other hand, log-parabolic energy spectra for relativistic electrons emitting synchrotron radiation can be generated through a statistical acceleration mechanism, where the probability of acceleration varies with energy (Massaro et al., 2006).
@@
In the spectral analysis, the model that best represents the observations can be inferred from the reduced chi-squared ($\chi^{2}_{r}$) values from the spectral fitting. It was found that, in general, the X-ray spectrum of the source is best represented by the BPL model (14 observations) and the log-parabola model (8 observations), suggesting that the simple PLmodel is not favored by the observations.
@@
We performed a spectral analysis of PKS 2155-304 using two approaches: a model-independent measure of the spectrum represented by the HR ratio and fitting the spectra with three spectral models commonly used in AGN studies. The HR analysis reveals that, although we observed hints of either a ’harder-when-brighter’ or ’softer-when-bright’ trend in some cases, most observations do not follow a simple trend in the HR-flux plane. In the context of AGN, the ’harder-when-brighter’ trend refers to a phenomenon observed in some blazars where the spectrum of emitted radiation shifts towards higher energies (becomes harder) as the brightness of the blazar increases. Conversely, the ’softer-when-bright’ trend occurs when the spectrum shifts towards lower energies (becomes softer) as the brightness increases. Such trends have been frequently observed across various frequency bands, e.g., optical, X-ray, and gamma-ray, as blazar flux fluctuates over time (see e. g., Das & Chatterjee, 2023; Noel et al., 2022; Bhatta et al., 2018; Bhatta & Webb, 2018).
@@
Observations of such definitive trends often point to a co-spatial population of particles contributing to the observed emission, which are seen in several blazars across various EM bands, Conversely, a complex trend in the HR-flux plane often indicates that the observed emission could be contributed by multiple independent emission regions. Alternatively, complex stochastic particle acceleration in turbulent jets, followed by cooling from emission, can also lead to a complex spectral profile.

### v2

@@ 4.3 Histograms and Log-normality
Similarly the expression for binormal distribution composed of two sub-populations having means $\mu_{1}$ and $\mu_{2}$ and standard deviations $\sigma_{1}$ and $\sigma_{2}$ is
@@
where $m_{1}$, $m_{2}$ and $s_{1}$, $s_{2}$ are location and scale parameters, respectively, for the two superposed lognormal distributions in the bimodal lognormal distribution. Also, the fit parameters obtained from fitting the normal and lognormal distributions are shown in Table 3.
@@
We found that seven observations (with IDs 0080940101, 0124930501 0158961301, 0411780301 and 0411780301) out of 22 observations exhibit a bimodal distribution. In Figure 4, we present the histogram for one of the observations, which shows a clear bimodal distribution. Blazar flux distributions showing bimodal feature are also reported in H. E. S. S. Collaboration et al. (2010); Mohorian et al. (2022); Shah et al. (2018). The presence of the bimodal distribution suggests a clear distinction between low and high flux states. However, since the light curves often exhibit continuous flux variation, it is difficult to categorize specific flux points as belonging to either a quiescent state or a flaring state using a bimodal PDF.
@@ 4.4 Power spectral density analysis
Studying variability in AGN using methods based on both the time domain and frequency domain can help uncover its underlying physical processes. The Power Spectral Density (PSD) serves as an important analytical tool in the frequency domain, measuring the ‘variability power’ at a given temporal frequency (or timescale). This enables us to understand the nature of variability and provides clues about its possible origin. The Discrete Fourier Transform (DFT) is widely used as a measure of PSD and is suitable for a discrete set of observations (see Vaughan et al., 2003; González-Martín & Vaughan, 2012; Mohorian et al., 2022). If events occur at discrete times $t_{j}$ with $j=1,2,...,n$, and sampled at frequencies $\nu_{\rm min}=\frac{1}{T}$, $2\nu_{\rm min}$ , $\nu_{\rm max}=\frac{1}{2\Delta t}$, DFT be expressed as,
@@
In this study, we performed PSD fit using the optimization algorithm for curve fitting included in the SciPy python library  [^3] https://docs.scipy.org/doc/scipy/reference/generated/scipy.optimize.curve_fit.html. Our analysis of the XMM-Newton X-ray light curves of PKS 2155-304 reveals a predominance of red noise, which is well characterized by a PL model. The PSD slopes ($\beta_{\mathrm{P}}$), as illustrated in Figure 6 and quantified in column 8 of Table 2, exhibit significant variability across the 22 observations of PKS 2155-304. The mean $\beta_{\mathrm{P}}$ value is 2.49, with a range from 0.922 (observation 0411780701) to 4.491 (observation 0411780201). These variable PSD slopes correspond to different stochastic processes: flicker noise ($\beta_{\mathrm{P}}\sim 1$), red noise ($\beta_{\mathrm{P}}\sim 2$), and other noise processes with steeper PL index, that is,   $\beta_{\mathrm{P}}>2$ (Press, 1978).
@@ 4.5.1 HARDNESS RATIO
It is interesting to note that although we observed pronounced flux variability throughout the observations, spectral variations measured by HR exhibit a complex pattern, with no significant variation in the majority of observations. Obs. IDs 0080940101, 0124930501, and 0411780501 reveal a trend (or a weak trend) of harder-when-brighter, whereas, Obs. ID 0411780201 shows a trend of softer-when-brighter (e. g., see Bhatta et al., 2018). In many observations, the error in HR at lower count rates is also higher, thus not providing conclusive HR variations with count rate. Additionally, Figure 8 presents the distribution of HR across the entire set of observations as a function of count rate. Note that different observational epochs are plotted in distinct colors. Although the data reveal no simple correlation between flux and HR, it is observed that long-term variability is characterized predominantly by achromatic variations, as indicated by the nearly flat behavior of HR distribution, suggesting that the spectral shape remains largely stable despite changes in overall flux. In contrast, chromatic variations, where the spectral shape undergoes considerable changes, are primarily observed on shorter timescales. Similar spectral behavior has been observed for a few sources in the optical band (see e.g., Raiteri et al., 2023; Otero-Santos et al., 2024).
@@ 4.5.2 SPECTRAL FITTING
We also performed the spectral analysis of the X-ray emission from blazar PKS 2155―304 by extracting the spectra of the observations within the energy range 0.3–10 keV and, consequently, fitted them with PL models from XSPEC [^4] https://heasarc.gsfc.nasa.gov/xanadu/xspec/manual/XspecManual.html. To accurately characterize the underlying spectral shape of the emission, we employed the following three different PL models.
(i) Power law (PL): This is a simple PL model in which the photon count rate distribution over the energy range is expressed solely by a single power index. The model can be written as
@@
(ii) Broken power law (BPL): In a BPL model, a spectral shape is characterized by two PL indices that meet at a discontinuity at an energy point known as the break energy. The distribution can be expressed as:
$\dfrac{dN}{dE}=\begin{cases}K\cdot E^{-\Gamma_{1}}&\text{if $E\geq E_{b}$},\\ K\cdot E^{-\Gamma_{2}}&\text{otherwise},\end{cases}$ (18)
Here $K$, $\Gamma_{1}$, $\Gamma_{2}$ and $E_{b}$ are normalization constant, low- and high-energy photon indexes and break energy, respectively.
(iii) Log parabola (LP): LP is one of the models widely used in the spectrum analysis of blazars (Massaro et al., 2004, 2006). The distribution function for LP is
@@
In this study, we fit the X-ray spectra obtained from XMM-Newton observations listed in Table 1 using three models available in the X-ray fitting package, xspec. To minimize instrumental artifacts, we restricted our analysis to the 0.3–10 keV energy range. We initially applied the simple PL model, which failed to yield a significant reduced chi-square ($\chi^{2}_{r}$). To improve the fit, we then applied more complex models: a BPL and LP. Both these models resulted in significantly improved $\chi^{2}_{r}$ values. Furthermore, we noted spectral curvature at lower energies. To accurately capture this feature, we utilized the LP model, a variant of the standard log-parabola model described in Equation 19. The results from our spectral fits are presented in Table 4, and a representative fitted spectrum from one observation is shown in Figure 9. The spectral fits for the remaining observations are presented in the Appendix. Our results indicate that the BPL and LP models each fit distinct subsets of observations better than the PL model, achieving better overall fits for the complete dataset and suggesting the presence of spectral curvature. A similar investigation of the spectral curvature of the PKS 2155-304 source was conducted by Gaur et al. (2017), utilizing XMM-Newton data from 2009 to 2014 and employing log-parabola model fitting techniques.
Table 4: X-ray spectral properties of the blazar PKS 2155-304. Col. 2: observation ID; Col. 3: Spectral models, power law (PL), broken power law (BPL); Col 4: photon index (PL), high-energy photon index (BPL); Col. 5: curvature parameter (LP), low-energy photon index (BPL); Col. 6: break/peak energy in keV; Col. 7: $\chi^{2}$/degrees of freedom; .
@@ 5.1 Timing Analysis
Variability: The source displays substantial variability both on intra-day and long-timescale. As indicated in Table 2, on intra-day timescales, it exhibits an average fractional variability of $\sim 6\%$, with one of the observations showing variability as high as 20%. During the entire observational period, the fractional variability reaches as high as $\sim 47\%$. In other words, the variability exhibits significant modulation with maximum flux levels approximately eight times higher than the minimum observed values.
The X-ray emissions of TeV blazars exhibit significant variability at intra-day timescales. This could possibly be attributed to the fact that the observed X-ray emission in HBLs is associated with synchrotron emission by a population of electrons located at the high-energy tail within the synchrotron component of their SEDs. However, it is also possible that LSPs and FSRQs also show intense X-ray variability. However, the number of these sources detected at TeV energies is much smaller, leading to a sample of TeV blazars that is heavily biased toward HSP blazars. This suggests that strong X-ray variability may be a general characteristic of the broader blazar population.
The IDV can be explained in the scenario of the turbulent jet of the blazar. In this framework, the observed variability can be interpreted as a superposition of fluctuations arising from numerous small-scale inhomogeneous components with varying Doppler boosting. Furthermore, as shocks propagate through the turbulent jet, energy dissipates from the largest to the smallest cells. Emission from some of these dominant smaller cells can contribute to short timescale variability. (Calafut & Wiita, 2015; Marscher, 2014; Bhatta et al., 2013). Using the variability timescale and the causality argument, the spatial scale of the inhomogeneity in the jet can be related to the size of the turbulent cells. An upper limit estimate for size of the compact spherical cells at redshift ($z$), corresponding to the minimum variability timescale ($\tau_{\rm var}$) and Doppler factor ($\delta$), can be computed using the following relation:
$R\leq\frac{\delta}{(1+z)}\cdot{c\tau_{\rm var}},$ (20)
where $\delta$ and $z$ represent bulk Doppler factor and redshift parameter, respectively. The lower limit of the Doppler factor associated with the source is of the order $\delta\sim$10 (Falomo et al., 2014; Kapanadze et al., 2020), which corresponds to a size of the emitting central region $R\sim 10^{14}$ cm using an average of minimum variability timescale over all the observations 2.0 ks. Similar results using Suzaku and NuSTAR observations are discussed in (Bhatta et al., 2018; Zhang et al., 2021). Furthermore, the variability timescales can be linked with the characteristics time scales such as cooling timescales of the source. In this particular case, the synchrotron cooling timescale in the observer’s frame is
@@
where $\gamma$ and $B$ represent bulk electron Lorentz factor and magnetic field in Gauss, respectively. An average cooling timescale of 2 ks, in combination with a magnetic field strength of B = 1 Gauss, typical for the jet’s magnetic field, results in an average electron energy of 4.32 × 104 Lorentz factors, enabling X-ray emission in the XMM-Newton band.
A linear RMS-flux relation is widely observed in AGNs and microquasars. This correlation is often linked to the multiplicative, non-linear stochastic processes driving the observed variability, which result in a flux distribution skewed toward higher flux levels, such as a lognormal PDF (Uttley et al., 2005). Conventionally, the linear RMS-flux relation, accompanied by a lognormal flux distribution, is attributed to variations in the viscosity parameter, $\alpha$, of the accretion disk. These fluctuations, driven by changes in viscosity at various radii, propagate outward and modulate mass accretion rates on larger scales (Lyubarskii, 1997). In the jet scenario, simulations of multi-frequency light curves demonstrate that this relationship can emerge when internal shocks of varying amplitudes occur. The shock strength is determined by the relative velocities of colliding plasma blobs within the jet (Kundu et al., 2022). However, as demonstrated in Figure 3, an analysis of the long-term observations reveals neither a linear RMS-flux relation nor any other definitive trend in variability behavior. This departure from a clear trend may be attributed to the complex interplay of multiple processes, including stochastic particle acceleration (e.g., Giannios et al., 2010; Marscher, 2014), radiation propagation through curved jet geometries (e.g., Raiteri et al., 2017; Bhatta, 2018; Young, 2010), and the intricate coupling between jet dynamics and the central engine (e.g., McKinney et al., 2012; Tchekhovskoy & Giannios, 2015).
The intra-day flux distributions of the blazar were analyzed using countrate distributions. Of the 22 observations, seven exhibited bimodal Gaussian behavior, while the remaining 15 showed unimodal Gaussian characteristics. The overall X-ray flux histogram revealed a complex structure with multiple peaks, suggesting that the distribution deviates from a standard PDF. In the context of long-term MWL variability studies, these findings can be compared with previous studies where optical and gamma-ray light curves demonstrated a strong correlation, and their respective flux distributions were best characterized by log-normal PDFs (Bhatta, 2021; Bhatta & Dhital, 2020). In the AGN literature, observations following log-normal PDFs are often interpreted as evidence that the observed variability arises from non-linear and multiplicative processes within the source.
On the other hand, the strong correlation between optical and gamma-ray emissions in blazars may be attributed to the fact that optical and Fermi/LAT gamma-ray observations are located near the synchrotron and high-energy peaks of the SED, respectively. However, X-ray emissions lie closer to the tail of the synchrotron component, probing the highest-energy electron population. The absence of a simple PDF in the X-ray regime suggests that the emission likely arises from multiple components and involves various emission processes within the source.
@@
Our analysis reveals that the observed PSD slopes ($\beta_{P}$) for PKS 2155-304 predominantly fall within the red noise regime $(2\leq\beta_{P}\leq 4)$, with occasional instances approaching pink noise $(\beta_{P}\approx 1)$. These variable PSD slopes might suggest distinct local manifestations of stochastic processes occurring within the turbulent jet structure. However, it is also possible that the long-term light curve is actually characterized by a single flatter PL PSD. In this case, the intra-day observations represent only a small segment observed across different flux states, resulting in apparent variations in the local PSD slopes. However, we did not estimate the long-term PSD by combining all the observations, as such a PSD would be affected by the large gaps between observations spanning years. This would result in an inaccurate characterization of the long-term X-ray variability in the source.
@@ 5.2 Spectral Analysis
In general, AGN spectra may be composed of various emission components, such as soft X-ray excess, neutral iron emission lines around 6.4 keV, and parts of either a synchrotron hump or a Compton hump. However, blazar spectral emission is primarily jet-dominated due to beamed electromagnetic radiation and the non-thermal emission in blazars usually follows a PL spectral profile (Aharonian et al., 2005b). Studying the X-ray spectra of blazars helps us understand the particle acceleration mechanisms at play, the properties of the jet environment, and the overall emission processes that power these luminous objects. The excellent sensitivity of XMM-Newton provides detailed information about the emission mechanisms, which we obtain by fitting the source X-ray spectra.
@@
A notable characteristic of our observations is the contrast between flux and spectral behavior. While pronounced flux variability was detected throughout the observing period, significant spectral variations, as measured by the HR, were observed in only two out of the total observations. The majority of observations exhibited no substantial spectral changes despite clear flux variations. To investigate this spectral behavior in greater detail, we performed comprehensive spectral analysis using three widely used XSPEC spectral models: PL, BPL, and LP. These models are chosen to represent non-thermal emission from relativistic electrons in the magnetic field of the blazar jet. In the synchrotron emission scenario, a simple PL spectrum can be interpreted as coherent synchrotron emission produced by a simple PL distribution of particles (Rybicki & Lightman, 1986). However, the BPL and LP models suggest more complex scenarios. A BPL might indicate that the observed emission is contributed by multiple populations of particles with different distributions. On the other hand, log-parabolic energy spectra for relativistic electrons emitting synchrotron radiation can be generated through a stochastic acceleration mechanism, where the probability of acceleration varies with energy (e. g., see Lemoine et al., 2024; Tramacere et al., 2007; Massaro et al., 2006).
@@
In the spectral analysis, the model that best represents the observations can be inferred from the reduced chi-squared ($\chi^{2}_{r}$) values from the spectral fitting. It was found that, in general, the X-ray spectra of the source were best represented by the BPL model (14 observations) and the LP model (8 observations), suggesting that the simple PL model is not favored by observations.
@@
We performed a spectral analysis of PKS 2155-304 using two approaches: a model-independent measure of the spectrum represented by the HR ratio and fitting the spectra with three spectral models commonly used in AGN studies. The HR analysis reveals that, although we observed hints of either a ‘harder-when-brighter’ or ‘softer-when-bright’ trend in some cases, most observations do not follow a simple trend in the HR-flux plane. In the context of AGN, the ‘harder-when-brighter’ trend refers to a phenomenon observed in some blazars where the spectrum of emitted radiation shifts towards higher energies (becomes harder) as the brightness of the blazar increases. Conversely, the ‘softer-when-bright’ trend occurs when the spectrum shifts towards lower energies (becomes softer) as the brightness increases. Such trends have been frequently observed across various frequency bands, e.g., optical, X-ray, and gamma-ray, as blazar flux fluctuates over time (see e. g., Das & Chatterjee, 2023; Noel et al., 2022; Bhatta et al., 2018; Bhatta & Webb, 2018).
The spectral behavior patterns, specifically the harder-when-brighter and softer-when-brighter trends, can be attributed to distinct physical processes in the jet. In the harder-when-brighter scenario, the correlation likely results from the acceleration of particles to higher energies during enhanced emission states, effectively extending the synchrotron spectrum to higher frequencies. This behavior suggests efficient particle acceleration mechanisms, possibly due to increased magnetic reconnection events or stronger shock acceleration. Conversely, the softer-when-brighter trend may indicate cooling dominance, where increased emission is accompanied by more efficient radiative cooling of the highest-energy electrons. This could occur when the particle injection rate increases while the maximum electron energy remains relatively constant, leading to a steepening of the electron energy distribution (see Mastichiadis & Kirk, 1997; Kirk et al., 1998; Böttcher, 2007, for details).
@@
Observations of such definitive trends often point to a co-spatial population of particles contributing to the observed emission, which are seen in several blazars across various EM bands, Conversely, a complex trend in the HR-flux plane often indicates that the observed emission could be contributed by multiple independent emission regions. Alternatively, complex stochastic particle acceleration in turbulent jets, followed by cooling from emission, can also lead to a complex spectral profile (see e. g., Marscher, 2014; Zhang et al., 2023).